In an emergency stop to avoid an accident, a shoulder strap seat belt holds a 51.5 kg passenger firmly in place. If the car were initially traveling at 85.0 km/h and came to a stop in 6.1 s along a straight, level road, what was the magnitude of the average force applied to the passenger by the seatbelt?
Answer #1

Solution :-
    Mass, m = 51.5 kg

    Initial velocity, U = 85 km/h
                             = 85 * ( 5 /18 ) m/s
                             = 23.61 m/s

    Final velocity, V = 0 m/s

    time , t = 6.1 s

    acceleration, a = ( V - U ) / t
                          = - 23.61/ 6.1
                          = - 3.87 m/s^2

    Force, F = m a
= (51.5)*( - 3.87)
= - 199.305 N        ( '-'sign represents retardation force )